This problem is for a project we were given:
Jeremy inherits $200,000 and invests part of the money in bonds with an interest rate of 8%. The remainder of the money is in savings at a 5% rate. What AMOUNT has he invested at EACH RATE if he recieves $14,800 interest for 1 year.?
So far I have the following:
let x= dollars invested at 8%
$200,000-x = dollars invested at 5%.
Equation as .08x+ .05(200,000-x)=$14,800
I know you have to clear the decimal and go from there, but when I tried to solve it I got some funky number.
thank you for your help!

Found 3 solutions by rajagopalan, checkley77, Earlsdon:
Answer by rajagopalan(174)

You have solved a major portion:
let x= dollars invested at 8%
$200,000-x = dollars invested at 5%.
Equation as .08x+ .05(200,000-x)=$14,800
From here 3 more steps..
1. Expand...
0.08x+10000-0.05x=14800
2. Simplify
0.03x=14800-10000=4800
3. Solve
0.03x=4800
x=4800/0.03=160000
So ..Answer
dollars invested at 8%=160000
dollars invested at 5%=40000
good day

Answer by checkley77(12844)

.08x+.05(200,000-x)=14,800
.08x+10,000-.05x=14,800
.03x=14,800-10,000
.03x=4,800
x=4,800/.03
x=$160,000 invested @ 8%
200,000-160,000=$40,000 invested @ 5%
Proof:
.08*160,000+.05*40,000=14,800
12,800+2,000=14,800
14,800=14,800

Answer by Earlsdon(6294)

Well, there's certainly nothing wrong with your initial equation!
0.08x + 0.05($200,000-x) = $14,800 Perform the indicated multiplication
0.08x+$10,000-0.05x = $14,800 Combine like-terms.
0.03x+$10,000 = $14,800 Subtract $10,000 from both sides.
0.03x = $4,800 Divide both sides by 0.03
x = $160,000 and...
$200,000-x = $200,000-$160,000 = $40,000
So $160,000 was invested at 8% while $40,000 was invested at 5%